# Neural Operator for Accelerating Coronal Magnetic Field Model

### v2

@@
Abstract.
Studying the sun’s outer atmosphere is challenging due to its complex magnetic fields impacting solar activities. Magnetohydrodynamics (MHD) simulations help model these interactions but are extremely time-consuming (usually on a scale of days). Our research applies the Fourier Neural Operator (FNO) to accelerate the coronal magnetic field modeling, specifically, the Bifrost MHD model. We apply Tensorized FNO (TFNO) to generate solutions from partial differential equations (PDEs) over a 3D domain efficiently. TFNO’s performance is compared with other deep learning methods, highlighting its accuracy and scalability. Physics analysis confirms that TFNO is reliable and capable of accelerating MHD simulations with high precision. This advancement improves efficiency in data handling, enhances predictive capabilities, and provides a better understanding of magnetic topologies.
Tensorized Fourier Neural Operator (TFNO), Magnetohydrodynamics (MHD), Nonlinear Force-Free Fields (NLFFF)
ccs: Applied computing Astronomyccs: Mathematics of computing Partial differential equationsccs: Computing methodologies Machine learning approaches
1. Introduction
@@
Studying the solar outer atmosphere has been a longstanding challenge in astrophysics, with its complex magnetic structures playing a crucial role in various solar phenomena (Wang et al., 2022; He et al., 2023; Borovsky et al., 2024). In the convection zone, the gas pressure dominates the magnetic field pressure, causing the plasma to move and carry the field along with it. These movements generate energetic flows and mass transfer from the chromosphere to the corona. Most of the energy conveyed to the outer solar atmosphere, as a result of exertion on the magnetic fields, is emitted in the chromosphere. In addition, it is in the chromosphere that the dynamics transition from being dominated by gas pressure to being dominated by magnetic force. Consequently, understanding and precisely simulating these intricate interactions are imperative for the progression of our understanding of space weather and its ramifications on the planet Earth.
@@
Coronal magnetic field modeling is mainly divided into magnetohydrodynamics (MHD) models (Braginskii, 1965; Yamasaki et al., 2022; Inoue et al., 2023), magnetohydrostatics (MHS) models (Shafranov, 1958), force-free models (Woltjer, 1958; Jing et al., 2010, 2014), and potential-field models (Wang and Sheeley Jr, 1992). These models solve specific partial differential equations (PDEs), ranging from complex magnetohydrodynamics processes to simplified assumptions of current-free conditions, to simulate and understand the structure and dynamics of the coronal magnetic field. These equations are discretized and solved on a grid, but the wide range of spatial and temporal scales and the complex nonlinear interactions between these scales require fine grids and high resolution, making these modeling extremely expensive. In recent years, there have been many attempts by researchers to simulate the solar atmosphere by computing codes for MHD. Leenaarts et al. (2007) simulated the solar atmosphere from the convection zone to the corona. Den (2016) presented a three-dimensional MHD simulation code designed to simulate space plasma phenomena with applications in modeling the solar surface and global solar wind structure. Carlsson et al. (2016) used the Bifrost MHD to provide the community with a realistic simulated magnetic field of the Sun’s outer atmosphere by elucidating the complex interactions between the magnetic field and the plasma. The Bifrost MHD are also used for performance evaluation of coronal field reconstruction using various approaches and tools (Fleishman et al., 2017, 2019). The Bifrost-based MHD model incorporates many physical processes and offers higher spatial and temporal resolution, focusing on the chromosphere and corona. This makes it well-suited for our study of predicting magnetic field topologies in corona.
However, MHD simulations are extremely time-consuming. To reduce the computational cost of MHD simulations, researchers have attempted to solve the MHD equations using artificial intelligence (AI)/machine learning (ML) methods to simulate the solar exosphere. AI/ML methods can solve the PDE equations several orders of magnitude faster compared to the traditional approach by providing a fast solver that approximates or augments the traditional approach (Berg and Nyström, 2018; Han et al., 2018; Long et al., 2018; Malek and Beidokhti, 2006). More and more researchers have been trying to use ML methods to solve MHD equations. For instance, Yang and Shen (2019) devised a method for globally mapping solar wind parameters at the source surface using an Artificial Neural Network (ANN) approach. This model efficiently tackles the MHD equations, employing ANN to adeptly forecast solar wind conditions. However, a notable limitation of ANN models, as highlighted in their research, is their inherent challenge in fully capturing the complex and nonlinear dynamics of solar wind acceleration and interactions within the corona. Similarly, Baty and Vigon (2024) presents a groundbreaking method to simulate solar coronal magnetic fields using Physics-Informed Neural Network (PINN) (Karniadakis et al., 2021; Gnanasambandam et al., 2023). This technique integrates the core partial differential equations of solar physics directly into the neural network, facilitating accurate modeling of MHD equilibria in the solar corona. This progress highlights the powerful application of AI/ML techniques to improve the modeling and predictive analysis of magnetic fields in the solar coronal environment. However, a significant limitation of PINNs is their inability to adapt to changes in system configurations or parameters without retraining. For instance, when solving the high-dimensional Navier-Stokes equations, each parameter change typically requires a new and specifically trained model. This requirement not only increases the computational cost but also restricts the method’s capability to efficiently handle and capture multi-scale phenomena.
Fourier neural operator (FNO) (Li et al., 2020b) has been introduced to solve this issue and improve the scalability of deep learning methods. The FNO represents a significant advancement over traditional physics-based modeling, as it is mesh-free, can transfer solutions between different grid geometries, and is significantly faster than conventional PDE solvers. In addition, FNO demonstrates computational efficiency surpassing standard vision transformers, and it effectively captures global dependencies compared to convolutional neural networks, graphical neural networks, and other deep learning methods. Despite the numerous advantages associated with FNOs, there is no documented application of FNOs to the solar coronal magnetic field modeling.
In this study, Tensorized Fourier Neural Operator (TFNO) (Kossaifi et al., 2023) is employed to accelerate coronal magnetic field modeling (i.e., the Bifrost-based MHD model). TFNO incorporates tensor decomposition (Kolda and Bader, 2009; Shen et al., 2022) to model the kernel operator, which is data efficient and highly parallelizable with reduced memory requirement and better generalization. We use data from Carlsson et al. (2016) to demonstrate the effectiveness and efficiency of TFNO by comparing it with state-of-the-art deep learning methods such as the Vision Transformer (Dosovitskiy et al., 2020), CNN-RNN (Wang et al., 2016), and CNN-LSTM (Hara et al., 2018). In addition to the evaluation metrics from an AI/ML perspective, evaluation from a physics perspective is conducted to verify that our prediction from TFNO is reliable.
2. Problem Setting
2.1. Bifrost-based MHD Model
@@
Bifrost is a flexible and massively parallel code described in detail in (Gudiksen et al., 2011). The Bifrost-based MHD model represents the realistic simulations of the magnetic solar outer atmosphere, with the 3D magnetic topologies on an enhanced area on the Sun. Bifrost has evolved from earlier numerical codes developed by Nordlund et al. (1995) and Galsgaard and Nordlund (1996), these codes share a common core. One of the key features of the Bifrost-based MHD model is its ability to simulate the multi-scale dynamics of solar phenomena, ranging from small-scale magnetic reconnection events to large-scale coronal mass ejections (CMEs). It incorporates realistic physics processes, such as radiative transfer, non-equilibrium ionization, and thermal conduction, to provide a comprehensive understanding of solar atmospheric phenomena. Specifically, a staggered mesh explicit code that solves the standard MHD PDEs on a Cartesian grid as follows
(1) $\frac{\partial\rho}{\partial t}$ $=-\nabla\cdot(\rho\mathbf{u})$
(2) $\frac{\partial(\rho\mathbf{u})}{\partial t}$ $=-\nabla\cdot(\rho\mathbf{u}\mathbf{u}-\tau)-\nabla P+\mathbf{J}\times\mathbf{B}+\rho\mathbf{g}$
(3) $\mu\mathbf{J}$ =×$=\nabla\times\mathbf{B}$
(4) $\mathbf{E}$ $=\eta\mathbf{J}-\mathbf{u}\times\mathbf{B}$
(5) $\frac{\partial\mathbf{B}}{\partial t}$ =×$=-\nabla\times\mathbf{E}$
(6) $\frac{\partial e}{\partial t}$ $=-\nabla\cdot(e\mathbf{u})-P\nabla\cdot\mathbf{u}+Q$
@@
where $\rho$, $\mathbf{u}$, $e$, $\mathbf{B}$ are the density, the velocity vector, the internal energy per unit volume, and the magnetic field intensity respectively. $\tau$, $P$, $\mathbf{J}$, $\mathbf{g}$, $\mu$, $\mathbf{E}$ and $\eta$ are the stress tensor, the gas pressure, the electric current density vector, the gravitational acceleration, the vacuum permeability, the electric field vector and the magnetic diffusivity respectively. The quantity $Q$ can contain many terms, depending on the individual experiment. It could for instance contain a term from the chosen Equation Of State (EOS), a term containing the effect of the Spitzer thermal conductivity, a term from radiative transfer, etc. The EOS needed to close this set of equations can be anything from a simple ideal gas EOS to a complex EOS including detailed microphysics (Gudiksen et al., 2011).
2.2. Tensorized Fourier Neural Operators
The idea of the Fourier Neural Operator (FNO) is motivated by Green’s function (Zhao et al., 2023). To solve a differential equation $Lu=f(\textbf{x})$ with an inhomogeneous term $f(\textbf{x})$ and a linear operator $L$, one would express the solution $u(\textbf{x})$ as an integral over the domain $D$ where the problem is defined:
@@ 2.2. Tensorized Fourier Neural Operators
where $\textbf{x},\textbf{y}\in\mathcal{D}\subseteq\mathbb{R}^{d}$ and $G$ is the Green’s function of the operator $L$. If there exists a function $g$ such that $G(\textbf{x},\textbf{y})=g(\textbf{x}-\textbf{y})$, Equation (7) reduces to a convolution operation, and $G$ serves as the convolution kernel. FNO learns the operator kernel by directly optimizing the kernel within the Fourier frequency domain using the Fast Fourier Transform (FFT), which has proved to be an efficient way to reduce the computational cost. For a parametric PDE, Green’s function also depends on the parameter. According to (Li et al., 2020b), the kernel integral operator is defined by
@@
where $v(\textbf{x})$ is the input function and the kernel integral operator $\mathcal{K}_{\theta}$ is parameterized by a neural architecture with learnable parameters $\theta$. In particular, $h_{i}$ denotes the latent representation generated by a Fourier layer in a sequence of Fourier layers indexed by $i$.
Figure 1. (a) The architecture of the fourier neural operators; (b) Fourier layer.
[…]
Example.
@@
For a better understanding of the structure of a Fourier layer, we visualize it in Figure 1. The input at the Fourier layer $i+1$, $h_{i}(\textbf{x})$ undergoes two transformations. $h_{i}(\textbf{x})$ is transformed to the Fourier space where a linearly transform $\mathcal{R}$ filters out the high-level Fourier modes, and finally the inverse Fourier transform $\mathcal{F}^{-1}$ converts back into the original space. Unlike (Li et al., 2020a) in which message passing is used as in a graph neural network, FNO leverages the convolution theorem, which states convolution is equivalent to point-wise multiplication in the Fourier space. Therefore, FNO initializes complex-valued learnable parameters directly in the Fourier space. Following the notations in (Li et al., 2020b), Fourier integral operator is defined by
(9) $\mathcal{K}_{\theta}(v)h_{i}=\mathcal{F}^{-1}(\mathcal{F}(\mathcal{K}_{\theta})\cdot\mathcal{F}(h_{i}))=\mathcal{F}^{-1}(R_{\phi}\mathcal{F}(h_{i}))$
@@
where $R_{\phi}$ is the Fourier transform of $\mathcal{K}$, i.e., $R_{\phi}=\mathcal{F}(\mathcal{K}_{\theta})$, and output is real-valued in the latent space. In the implementation, the fast Fourier transform and its inverse is used to preserve computational efficiency. The bottom path input $h_{i}(x)$ applies a local linear transform $\mathcal{W}$, where the discretized $\mathcal{W}$ is a matrix of learnable parameters. The results of the two paths are then added together, followed by a nonlinear activation function $\sigma$.
With the Fourier layer defined, Figure 1 also shows the architecture of FNO. FNO incorporates a lifting layer $\mathcal{P}$ which lifts the input tensor into high-dimensional space. A sequence of Fourier layers follows to update the latent representations by
(10) $h_{i+1}=\sigma(\mathcal{W}_{i+1}h_{i}(\textbf{x})+\mathcal{K}_{\theta_{i+1}}(v)h_{i})(\textbf{x})$
with $i=1,2,\dots,N-1$. Since $R_{\phi}$, $\mathcal{W}$ differ in each layer, they are indexed by $i$ as well. Finally, layer $\mathcal{Q}$ projects $h_{N}$ to the solution function space denoted as $u(\textbf{x})$. In our case, the output is multi-channel images.
As FNO learns the operator to map from input space $v$ to the solution space $u$, it solves PDEs with variable parameters (Li et al., 2020b). For comparison, frameworks such as PINN (Raissi et al., 2019) only solve one specific setting of the PDEs by fixing the boundary/initial conditions and coefficients in PDEs. TFNO extends FNO by parameterizing the learnable weights $\mathcal{W}$ within the Fourier domain with a low-rank tensor factorization (e.g., Tucker, Tensor Train). Thus, TFNO imposes a low-rank constraint on the Fourier domain representation of the convolutional weights within the FNO. This factorization leads to a significant reduction in the number of parameters required to represent the operator, enhancing the model’s data efficiency and generalization capabilities (Kossaifi et al., 2023).
3. Experiments

### v3

# Global-local Fourier Neural Operator for Accelerating Coronal Magnetic Field Model

@@
Abstract
Exploring the outer atmosphere of the sun has remained a significant bottleneck in astrophysics, given the intricate magnetic formations that significantly influence diverse solar events. Magnetohydrodynamics (MHD) simulations allow us to model the complex interactions between the sun’s plasma, magnetic fields, and the surrounding environment. However, MHD simulation is extremely time-consuming, taking days or weeks for simulation. The goal of this study is to accelerate coronal magnetic field simulation using deep learning, specifically, the Fourier Neural Operator (FNO). FNO has been proven to be an ideal tool for scientific computing and discovery in the literature. In this paper, we proposed a global-local Fourier Neural Operator (GL-FNO) that contains two branches of FNOs: the global FNO branch takes downsampled input to reconstruct global features while the local FNO branch takes original resolution input to capture fine details. The performance of the GL-FNO is compared with state-of-the-art deep learning methods, including FNO, U-NO, U-FNO, Vision Transformer, CNN-RNN, and CNN-LSTM, to demonstrate its accuracy, computational efficiency, and scalability. Furthermore, physics analysis from domain experts is also performed to demonstrate the reliability of GL-FNO. The results demonstrate that GL-FNO not only accelerates the MHD simulation (a few seconds for prediction, more than ×20,000 speed up) but also provides reliable prediction capabilities, thus greatly contributing to the understanding of space weather dynamics. Our code implementation is available at https://github.com/Yutao-0718/GL-FNO.
Index Terms:
Global-local Fourier Neural Operator (GL-FNO), Coronal Magnetic Field Model, Deep Learning.
I Introduction
@@
Studying the solar outer atmosphere has been a longstanding challenge in astrophysics, with its complex magnetic structures playing a crucial role in various solar phenomena [1, 2, 3]. In the convection zone, the gas pressure dominates the magnetic field pressure, causing the plasma to move and carry the field along with it. These movements generate energetic flows and mass transfer from the chromosphere to the corona. As a result of exertion on the magnetic fields, most of the energy conveyed to the outer solar atmosphere is emitted in the chromosphere. In addition, in the chromosphere, the dynamics transition from being dominated by gas pressure to being dominated by magnetic force. Consequently, understanding and precisely simulating these intricate interactions are important for the progression of our understanding of space weather and its impact on the planet Earth.
@@
Coronal magnetic field modeling is mainly divided into magnetohydrodynamics (MHD) models [4, 5, 6, 7], magnetohydrostatics (MHS) models [8, 9], force-free models [10, 11, 12], and potential-field models [13]. These models solve specific partial differential equations (PDEs), ranging from complex magnetohydrodynamics processes to simplified assumptions of current-free conditions, to simulate and understand the structure and dynamics of the coronal magnetic field. These equations are discretized and solved on a grid, but the wide range of spatial and temporal scales and the complex nonlinear interactions between these scales require fine grids and high resolution, making these modeling extremely expensive.
In recent years, there have been many attempts by researchers to simulate the solar atmosphere by MHD. Leenaarts et al. [14] simulated the solar atmosphere from the convection zone to the corona. Den [15] presented a three-dimensional MHD simulation code designed to simulate space plasma phenomena with applications in modeling the solar surface and global solar wind structure. Carlsson et al. [16] used the Bifrost MHD to provide the community with a realistic simulated magnetic field of the Sun’s outer atmosphere by elucidating the complex interactions between the magnetic field and the plasma. The Bifrost MHD is also used for performance evaluation of coronal field reconstruction using various approaches and tools [17, 18]. The Bifrost-based MHD model incorporates many physical processes and offers higher spatial and temporal resolution, focusing on the chromosphere and corona. This makes it well-suited for our study of predicting magnetic field topologies in corona.
Coronal magnetic field modeling is extremely time-consuming. To reduce the computational cost, researchers have attempted to simulate the solar outer atmosphere using artificial intelligence (AI)/machine learning (ML) methods. AI/ML has been proven to be efficient in various applications [19, 20, 21, 22, 23, 24, 25]. Researchers have used AI/ML methods for coronal magnetic field modeling, especially using Physics-Informed Neural Network (PINN) [26, 27, 28]. PINN incorporates physical laws, typically in the form of PDEs, directly into the neural network’s loss function. The network is trained to minimize the residuals of the PDE and satisfy any boundary and initial conditions. For instance, Jarolim et al. [29] presented an approach for coronal magnetic-field extrapolation, using PINN that integrates observational data and the physical force-free magnetic-field model. The method flexibly finds a trade-off between the observation and force-free magnetic-field assumption, improving the understanding of the connection between the observation and the underlying physics. To further relax the force-free assumption for the dense plasma conditions in the lower atmosphere, Jarolim et al. [30] utilized multi-height magnetic field measurements in combination with PINN to advance solar magnetic field extrapolations. In addition, Baty and Vigon[31] explored the functionality of PINNs through application to two particular solar problems. First, they consider the computation of two-dimensional (2D) force-free magnetic equilibria representative of arcades and loop-like structures in the solar corona by solving an associated Grad–Shafranov-like equation. Second, their method is extended to a more complex system of differential equations that is an incompressible resistive MHD set, to compute 2D magnetic reconnection solutions.
Despite a significant advancement over traditional physics-based modeling, PINNs have the following issues (1) Limited Generalization: PINNs are designed to solve a specific PDE or a specific instance of a problem. They are well-suited for problems where the PDE is fixed including boundary/initial conditions. For a different PDE, a different PINN needs to be trained for the solution. (2) Training Scalability: Compared to standard neural networks, PINNs often require more complex and time-consuming training processes because they need to satisfy both data-driven and physics-based loss functions [32].
To address the above issues of PINNs, Fourier neural operator (FNO) [33, 34] has been introduced to learn a mapping between function spaces, effectively learning the solution operator for a family of PDEs. Instead of solving a single instance of a PDE, they generalize across different input conditions [35]. FNO is trained on a dataset of solutions to the PDE across different parameter settings or initial conditions. Once trained, they can quickly predict solutions for a new instance. A key reason for FNO’s success is its ability to accurately model long-range dependencies in spatiotemporal data by learning global convolutions in a computationally efficient manner [36]. Up to now, there is only one paper using FNO to learn MHD solar wind plasma parameters as output using the solar photospheric magnetic field observation as input [37]. There is no documented application of FNOs to directly derive the solution of coronal magnetic field modeling (i.e., the Bifrost-based MHD model).
In this paper, we propose a global-local Fourier Neural Operator (GL-FNO) to accelerate coronal magnetic field modeling (i.e., the Bifrost-based MHD model). GL-FNO contains two branches of FNOs: global FNO takes downsampled input to reconstruct global features while the local FNO takes original resolution input to capture fine details. In our global and local FNOs implementation, we incorporate tucker decomposition [38, 39] to model the kernel operator with reduced memory requirement and better generalization. To summarize, the contributions of this paper are
  • We proposed a novel method GL-FNO to obtain solutions for coronal magnetic field modeling (i.e., the Bifrost-based MHD model), enabling global and local features learning simultaneously.
  • We use data from [16] to demonstrate the effectiveness and efficiency of GL-FNO by comparing it with state-of-the-art deep learning methods such as FNO [33], U-NO [40], U-FNO [41], Vision Transformer [42], CNN-RNN [43], and CNN-LSTM [44]. Specifically, our GL-FNO achieves the best test accuracy compared to other deep learning methods.
  • In addition to the evaluation metrics from an AI/ML perspective, evaluation from a physics perspective is conducted to verify that our prediction from GL-FNO is reliable.
The rest of this paper is organized as follows. The Bifrost-based MHD Model is introduced in Section II, followed by our proposed GL-FNO in Section III. Our experiments and results are summarized in Section IV. Finally, conclusions are discussed in Section V.
Figure 1: (a) The architecture of the global-local fourier neural operators; (b) local fourier layer; (c) tucker decomposition.
II Bifrost-based MHD Model
@@
Bifrost is a flexible and massively parallel code described in detail in [45]. The Bifrost-based MHD model represents the realistic simulations of the magnetic solar outer atmosphere, with the 3D magnetic topologies on an enhanced area on the Sun. Bifrost has evolved from earlier numerical codes developed by [46] and [47], these codes share a common core. One of the key features of the Bifrost-based MHD model is its ability to simulate the multi-scale dynamics of solar phenomena, ranging from small-scale magnetic reconnection events to large-scale coronal mass ejections (CMEs). It incorporates realistic physics processes, such as radiative transfer, non-equilibrium ionization, and thermal conduction, to provide a comprehensive understanding of solar atmospheric phenomena. Specifically, a staggered mesh explicit code that solves the standard MHD PDEs on a Cartesian grid as follows
$\frac{\partial\rho}{\partial t}$ $=-\nabla\cdot(\rho\mathbf{u})$ (1)
$\frac{\partial(\rho\mathbf{u})}{\partial t}$ $=-\nabla\cdot(\rho\mathbf{u}\mathbf{u}-\tau)-\nabla P+\mathbf{J}\times\mathbf{B}+\rho\mathbf{g}$ (2)
$\mu\mathbf{J}$ =×$=\nabla\times\mathbf{B}$ (3)
$\mathbf{E}$ $=\eta\mathbf{J}-\mathbf{u}\times\mathbf{B}$ (4)
$\frac{\partial\mathbf{B}}{\partial t}$ =×$=-\nabla\times\mathbf{E}$ (5)
$\frac{\partial e}{\partial t}$ $=-\nabla\cdot(e\mathbf{u})-P\nabla\cdot\mathbf{u}+Q$ (6)
@@
where $\rho$, $\mathbf{u}$, $e$, $\mathbf{B}$ are the density, the velocity vector, the internal energy per unit volume, and the magnetic field intensity respectively. $\tau$, $P$, $\mathbf{J}$, $\mathbf{g}$, $\mu$, $\mathbf{E}$ and $\eta$ are the stress tensor, the gas pressure, the electric current density vector, the gravitational acceleration, the vacuum permeability, the electric field vector and the magnetic diffusivity respectively. The quantity $Q$ can contain many terms, depending on the individual experiment. It could for instance contain a term from the chosen Equation Of State (EOS), a term containing the effect of the Spitzer thermal conductivity, a term from radiative transfer, etc. The EOS needed to close this set of equations can be anything from a simple ideal gas EOS to a complex EOS including detailed microphysics [45].
The process of modeling the outer solar atmosphere is explained in [45]. Consider the example of the MHD PDE equation for the magnetic field intensity $\mathbf{B}$. Combining (4) and (5) results in the MHD Magnetic Induction Equation:
$\frac{\partial\mathbf{B}}{\partial t}$ =×$=\nabla\times(\mathbf{u}\times\mathbf{B})-\nabla\times(\eta\mathbf{J}).$ (7)
To solve this equation, Bifrost employs a staggered grid discretization approach using the finite difference method [48], which assigns the different components of the magnetic field and velocity to different locations within the grid cells. Subsequently, the curl of the magnetic field is computed using a sixth-order exact scheme to handle spatial derivatives. The velocity field and magnetic field interaction term $\mathbf{u}\times\mathbf{B}$ is interpolated to the desired grid location, after which the curvature is calculated. Bifrost employs either a third-order Runge-Kutta method or a third-order Hyman method for the calculations to advance the solution of the differential equations in time. Finally, Bifrost implements non-periodic boundary conditions using “ghost zones” [45]. Depending on the specific problem, different types of boundary conditions (e.g. symmetric, asymmetric, or extrapolated boundary values) are applied to these ghost zones. After performing these calculations, Bifrost can obtain the solution for the magnetic field $\mathbf{B}$.
III Proposed Method
Our proposed GL-FNO is a modification of FNO. In GL-FNO, we have a global iterative architecture: $h^{G}_{0}\mapsto h^{G}_{1}\mapsto\dots\mapsto h^{G}_{N}$ and a local iterative architecture: $h^{L}_{0}\mapsto h^{L}_{1}\mapsto\dots\mapsto h^{L}_{N}$, where $\{h^{G}_{i}\}_{i=0}^{N},\{h^{L}_{i}\}_{i=0}^{N}$ are a sequence of global and local functions, respectively. As shown in Fig. 1(a), the input $u(\textbf{x})$ is first lifted to a higher dimensional global representation $h^{G}_{0}=\mathcal{P}^{G}(u(\textbf{x}))$ and local representation $h^{L}_{0}=\mathcal{P}^{L}(u(\textbf{x}))$ by the transformations $\mathcal{P}^{G}$ and $\mathcal{P}^{L}$, which are usually parameterized by shallow fully-connected neural networks. Then we apply several iterations of global update $h^{G}_{i}\mapsto h^{G}_{i+1}$ and local update $h^{L}_{i}\mapsto h^{L}_{i+1}$ (introduced later). The output $v(\textbf{x})=\mathcal{Q}^{G}(h^{G}_{N}(\textbf{x}))+\mathcal{Q}^{L}(h^{L}_{N}(\textbf{x}))$ is the summation of the projections by the transformations $\mathcal{Q}^{G}$ and $\mathcal{Q}^{L}$.
Each iterative update is enabled by Fourier layers as
@@ III Proposed Method
where $\sigma$ is a nonlinear activation function, $\mathcal{K}^{G}_{\theta^{G}_{i+1}}$ or $\mathcal{K}^{L}_{\theta^{L}_{i+1}}$ is the kernel integral operator parameterized by a neural architecture with learnable parameters $\theta^{G}_{i+1}$ or $\theta^{L}_{i+1}$.
@@
For a better understanding of the structure of a Fourier layer, we visualize the Fourier layer of the local branch in Fig. 1(b). The input at the Fourier layer $i+1$, i.e., $h_{i}^{L}(\textbf{x})$, undergoes two transformations. In the top path, $h_{i}^{L}(\textbf{x})$ is transformed to the Fourier space through Fourier transform $\mathcal{F}$ followed by a linearly transform $\mathcal{R}$ filters out the high-level Fourier modes, and finally the inverse Fourier transform $\mathcal{F}^{-1}$ converts back into the original space. Unlike [49] in which message passing is used as in a graph neural network, the Fourier layer leverages the convolution theorem, which states convolution is equivalent to point-wise multiplication in the Fourier space. Therefore, the Fourier layer initializes complex-valued learnable parameters directly in the Fourier space. Following the notations in [33], Fourier integral operator is defined by
$\mathcal{K}^{L}_{\theta}h^{L}_{i}=\mathcal{F}^{-1}\left(\mathcal{F}(\mathcal{K}^{L}_{\theta}\right)\cdot\mathcal{F}(h_{i}^{L}))=\mathcal{F}^{-1}\left(R_{\phi}\mathcal{F}(h^{L}_{i})\right)$ (9)
@@
where $R_{\phi}$ is the Fourier transform of $\mathcal{K}^{L}_{\theta}$, i.e., $R_{\phi}=\mathcal{F}(\mathcal{K}^{L}_{\theta})$, and output is real-valued in the latent space. In the implementation, the fast Fourier transform and its inverse is used to preserve computational efficiency. The bottom path applies $h^{L}_{i}(x)$ with a linear transform $\mathcal{W}^{L}_{i+1}$, where the discretized $\mathcal{W}^{L}_{i+1}$ is a matrix of learnable parameters. The results of the two paths are then added together, followed by a nonlinear activation function $\sigma$.
In GL-FNO, we also parameterize the learnable weights within the Fourier domain with a low-rank tucker factorization as shown in Fig. 1(c). This factorization leads to a significant reduction in the number of parameters required to represent the operator, enhancing the model’s data efficiency and generalization capabilities [50].
Our GL-FNO learns the operator to map from input space $u$ to the solution space $v$ with global and local branches. The global branch captures the overall structure and dynamics, while the local branch focuses on fine-grained details. This fusion enables the model to generalize better across different physical conditions while retaining accuracy in areas requiring detailed resolution. Additionally, by integrating both scales, the network can learn complex interactions more efficiently, leading to faster and more accurate prediction in diverse physical systems.
IV Experiments and Results